On January 8, the end of the first weekly pay period of the year, Regis Company's payroll register showed that its employees earned $27,760 of office salaries and $60,840 of sales salaries. Withholdings from the employees' salaries include FICA Social Security taxes at the rate of 6.20%, FICA Medicare taxes at the rate of 1.45%, $13,560 of federal income taxes, $1,390 of medical insurance deductions, and $920 of union dues. No employee earned more than $7,000 in this first period.

Required:
1.1 Calculate below the amounts for each of these four taxes of Regis Company. Regis’s merit rating reduces its state unemployment tax rate to 5% of the first $7,000 paid each employee. The federal unemployment tax rate is 0.60%. (Round your answers to 2 decimal places.)

Regis Companys: Tax FICA-Social Security FICA-Medicare FUTA SUTA January 8 earnings subject to tax Tax Rate Tax Amount

1.2 Prepare the journal entry to record Regis Company's January 8 (employee) payroll expenses and liabilities. (Round your answers to 2 decimal places.)

Journal entry worksheet Record the employee payroll expenses and liabilities for the first weekly pay period of the year. Note: Enter debits before credits. Date General Journal Debit Credit Jan 08 Record entry Clear entry View general journal

2. Prepare the journal entry to record Regis’s (employer) payroll taxes resulting from the January 8 payroll. Regis’s merit rating reduces its state unemployment tax rate to 5% of the first $7,000 paid each employee. The federal unemployment tax rate is 0.60%. (Round your answers to 2 decimal places.)

Journal entry worksheet Record the employers payroll expenses and liabilities for the first weekly pay period of the year. Note: Enter debits before credits. Date General Journal Debit Credit Jan 08 Record entry Clear entry View general journal
Answer #1

Answer

Regis Company’s:

Tax

January 8 earnings subject to tax $

Tax Rate

Tax Amount $

FICA-Social Security

88600

6.20%

5493.20

FICA-Medicare

88600

1.45%

1284.70

FUTA

88600

0.60%

531.60

SUTA

88600

5.00%

4430.00

Office salaries= $27760

Sales salaries= $60840

Total of office and sales salaries adds up to $88600 i.e. total earnings.

By applying the tax rates to the total earnings we get tax amount in the last column.

88600*6.2%=5493.20

88600*1.45%=1284.7

88600*0.60%=531.6

88600*5%=4430

Journal entry to record Regis company's January 8 payroll expenses and liabilities.

Date

     General Journal

Debit

Credit

Jan 08

Office salaries expense

27760.00

Sales salaries expense

60840.00

FICA—Social sec. taxes payable

5493.20

FICA—Medicare taxes payable

1284.70

Employee fed. income taxes payable

13560.00

Employee medical insurance payable

1390.00

Employee union dues payable

920.00

Salaries payable

65952.10

(Recording of payroll expenses and liabilities)

Salaries payable is the balance remaining after deducting taxes, insurance and union dues from the total earnings.

(2)Journal entry to record Regis company's January 8 payroll expenses

Date

General Journal

Debit

Credit

Jan 08

Payroll taxes expense

11739.50

FICA—Social sec. taxes payable

5493.20

FICA—Medicare taxes payable

1284.70

State unemployment taxes payable

4430.00

Federal unemployment taxes payable

531.60

(Recording of payroll expenses)